# Gradient-flow adaptive importance sampling for Bayesian leave one out cross-validation for sigmoidal classification models

### v1

@@ Abstract
We introduce a set of gradient-flow-guided adaptive importance sampling (IS) transformations to stabilize Monte-Carlo approximations of point-wise leave one out cross-validated (LOO) predictions for Bayesian classification models. One can leverage this methodology for assessing model generalizability by for instance computing a LOO analogue to the AIC or computing LOO ROC/PRC curves and derived metrics like the AUROC and AUPRC. By the calculus of variations and gradient flow, we derive two simple nonlinear single-step transformations that utilize gradient information to shift a model’s pre-trained full-data posterior closer to the target LOO posterior predictive distributions. In doing so, the transformations stabilize importance weights. Because the transformations involve the gradient of the likelihood function, the resulting Monte Carlo integral depends on Jacobian determinants with respect to the model Hessian. We derive closed-form exact formulae for these Jacobian determinants in the cases of logistic regression and shallow ReLU-activated artificial neural networks, and provide a simple approximation that sidesteps the need to compute full Hessian matrices and their spectra. We test the methodology on an $n\ll p$ dataset that is known to produce unstable LOO IS weights.
@@ 1 Introduction
In Bayesian formulations of classification problems where one uses a vector of covariates $\mathbf{x}\in\mathbb{R}^{p}$ to estimate the probability of an outcome labeled by $y\in\{0,1\}$, one specifies a likelihood function $\ell$, where
where $\pi(\boldsymbol{\theta})$ is the joint prior distribution over the parameter vector $\boldsymbol{\theta}$, corresponding to the likelihood function $\ell(\boldsymbol{\theta})$. There are usually many plausible choices for $\ell,$ even within the same family of models (for example different sets of features in logistic regression) – one needs methodology to evaluate different models to inform decisions such as model selection.
One usually wishes to evaluate how a model will generalize to new data, by estimating its accuracy on data that is both unused during training and statistically independent from the training data. Simulating performance under unseen data requires either fitting the model multiple times (for k-fold cross validation, which can be expensive on large scale datasets, with large-scale models), or removing a subset of the data entirely (in the case of test/train splits).
Other general methods of assessing model generalizability exist, notably information criterion related to the Akaike Information criteria (AIC). The AIC and Bayesian variants [Gelman et al., 2014, Watanabe, 2013, Vehtari et al., 2017, Piironen and Vehtari, 2017a] are asymptotic approximations [Stone, 1977, Watanabe, 2010] of a model’s leave one out (LOO) cross validation loss. In their standard usage, these methods approximate the total log LOO likelihood for a given model/dataset, a quantity that does not directly map to ROC/PRC curves. However, similar methods can be used to compute LOO versions of these curve. The receiver operator curve (AUROC) [Hanley and McNeil, 1982] and the area under the precision recall cure (AUPRC) are familiar, interpretable, and often desirable for use in decision making processes.
In this manuscript, we motivate an adaptive importance sampling method for LOO, based on transformations derived by reduction of a given functional criterion by stepping down a gradient flow. These transformations use model-specific gradient information – we derive explicit formulae for the transformations and their Jacobian determinants in the case of common sigmoidal classification models (logistic regression and Bayesian ReLU-activated neural networks).
@@ 2 Preliminaries
2.1 Notation
We denote vectors using bold-faced lower case symbols, and matrices using bold-faced uppercase symbols. By convention, all vectors are assumed to be column vectors unless otherwise stated. Given a matrix $\mathbf{W}=(w_{ij})$, the $i$-th row is denoted $\mathbf{w}_{i,}$ and $j$-th column is denoted $\mathbf{w}_{,j}$.
@@ 2.1 Notation
2.2 Importance sampling-based approximate leave one out cross validation (IS-LOO)
Suppose that one has pre-trained a Bayesian classification model such that one is able to sample its posterior parameters $\boldsymbol{\theta}_{s}\overset{\mathrm{iid}}{\sim}\pi(\boldsymbol{\theta}|\mathcal{D}).$ Our objective is to use knowledge of this full-data posterior distribution to estimate how the model would behave if any single point is left out at training. One can relate the full-data model to the model with observation $i$ left out using the Bayesian update equation
@@ 2.2 Importance sampling-based approximate leave one out cross validation (IS-LOO)
which is a Fredholm second-kind integral equation with respect to $\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)}).$ This integral equation is in-practice difficult to solve due to the typically-high dimensionality of $\boldsymbol{\theta}.$
Rather than attempt direct inversion of Eq. 3, our starting point is the observation that
@@
which provides the ratio of densities between a distribution we know (the full-data posterior $\pi(\boldsymbol{\theta}|\mathcal{D})$) and a distribution whose statistics we would like to compute (the point-wise LOO posterior $\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})$). To use the former in order to compute statistics of the latter we turn to importance sampling, noting that for an integrable function $f,$
@@
$=\int f(\boldsymbol{\theta})\frac{\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})}{\pi(\boldsymbol{\theta}|\mathcal{D})}\pi(\boldsymbol{\theta}|\mathcal{D})\mathrm{d}\boldsymbol{\theta}$
$=\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}}\left[f(\boldsymbol{\theta})\nu_{i}(\boldsymbol{\theta})\right].$ (5)
Sampling $\boldsymbol{\theta}_{k}\overset{\mathrm{iid}}{\sim}\pi(\boldsymbol{\theta}|\mathcal{D})$, we can approximate Eq. 5 using the Monte-Carlo integral
@@
2.3 LOO-based metrics
In Bayesian workflows, it is common to evaluate model generalizability by using importance sampling LOO to compute a Bayesian analogue to the Aikaike Information Criterion (AIC). The LOO information criterion (LOO-IC) takes the form
@@ 2.3 LOO-based metrics
2.4 Weight stabilization
Often it is the case that using the computed posterior $\pi(\boldsymbol{\theta}|\mathcal{D})$ as the proposal distribution for importance sampling has slow convergence properties – the $1/\ell$ importance weights are known to have large or unbounded variance [Peruggia, 1997], making the importance sampler estimate for LOO noisy.
Two practical model agnostic methods for controlling the tail of the importance weights are through weight truncation [Ionides, 2008] and Pareto smoothing  [Vehtari et al., 2015, 2017]. Pareto smoothing replaces the largest $M$ weights with their corresponding rank-values from a fitted generalized Pareto-distribution [Zhang and Stephens, 2009]. Pareto smoothed importance sample (PSIS)-based LOO implementations are widely available in software packages such as Stan and ArviZ. However, PSIS-LOO is insufficient in problems where the Pareto distribution does not well-fit the tail distribution of importance weights – where the estimated Pareto shape parameter $\hat{k}$ exceeds $0.7$ – this parameter is a useful criteria for accessing the validity of a given importance sampler. In such a case, it would be desirable to perform an additional model-specific controlled transformation on the proposal distribution that will induce more efficient computations.
@@ 2.5 Adaptive importance sampling
The objective of adaptive importance sampling [Bugallo et al., 2017, Cornuet et al., 2011, Elvira and Martino, 2022] in the context of LOO is to use a transformation to nudge, independently for each observation $i,$ the posterior distribution closer to the LOO distribution $\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})$ (relationships between the different distributions is depicted in Fig. 1). First we’ll examine importance sampling under an arbitrary bijective transformation.
@@
2.6 Adaptive importance sampling using variational posteriors
For computational expediency, variational methods are often used in place of MCMC for Bayesian inference, obtaining a variational approximation $\hat{\pi}(\boldsymbol{\theta}|\mathcal{D})$ to the true posterior, where $\hat{\pi}$ lies within a given family of transformed probability distributions. In problems where one expects a substantial discrepancy between the true posterior and $\hat{\pi},$ one may correct for this discrepancy by noting that
@@ 3 Methods
Eq. 13 is valid for an arbitrary bijection $T_{i}.$ The objective in using transformations is to shift the proposal distribution closer to the targeted LOO distribution for each observation – to partially invert Bayes rule. First, we motivate three different single-step transformations of the form
@@
3.1 Single step transformations
@@
Starting from the associated variational problem (Appendix S.1.2), and applying the same rationale that went into developing the KL-descending transformation, one arrives at the single step variance-reducing transformation,
@@
Log-likelihood descent: The two single step transformations of Eq. 21 require computing both the gradient and Hessian of the trained full-data posterior. We also consider a simpler alternate transformation [Elvira et al., 2023]
$T^{\textrm{LL}}_{i}(\boldsymbol{\theta})$ $=\boldsymbol{\theta}-h\underbrace{\nabla\log\ell(\boldsymbol{\theta}|y_{i},\mathbf{x}_{i})}_{-Q_{i}^{\textrm{LL}}},$ (23)
equivalent to repelling each parameter sample a step $h>0$ away from the maximization of the log likelihood function for point $i.$
3.2 Resolving the posterior density
Both the KL (Eq. 21) and variance (Eq. 22) descent transformations take steps proportional to the posterior density $\pi(\boldsymbol{\theta}|\mathcal{D}).$ If a variational approximation for $\pi({\boldsymbol{\theta}}|\mathcal{D})$ is available, using it in Eqs. 21 and 22 as a stand-in for the posterior density helps simplify the computation of the transformations and their Jacobians, particularly when using mean-field or low-order ADVI.
In the absence of variational approximation, one may evaluate the posterior densities exactly using Bayes rule, absorbing the unknown normalization constant $\mathcal{Z}$ into the step size $h.$ Then, the KL and variance descent transformations are
@@ 3.2 Resolving the posterior density
and
@@
3.3 Step size selection
The KL-divergence and variance descent transformations correspond to a forward Euler solver on the respective gradient flow equations. According to linear stability analysis, Euler’s method has the conditional stability criteria $h<{2}/{\max_{k}|\mathrm{Re}(\lambda_{k})|}$ where $\lambda_{k}$ are the eigenvalues of the Jacobian of the system (Jacobians of the functions $Q_{i}$). In each case the structure of the Jacobian admits cheap approximations of $\lambda_{k}$. However, for nonlinear systems, this criterion is not sufficient for achieving stability.
@@ 3.3 Step size selection
$h_{i}=\rho\min_{s,\alpha}\left\{\left|\frac{\sqrt{\Sigma_{\alpha,\alpha}}}{Q_{i}(\boldsymbol{\theta}_{s})_{\alpha}}\right|\right\}$ (27)
@@
3.4 Jacobian determinant approximation
For any of the three single-step transformations, one may approximate $\left|\mathbf{J}_{T_{i}}\right|$ by noting that
@@ 3.4 Jacobian determinant approximation
and truncating to $\mathcal{O}(h),$ sidestepping the computation of Hessian matrices and their spectra. Note that any higher order terms in this expansion require characterization of the spectra of $\nabla Q_{i},$ for each observation $i$, and for each sampled parameter $\boldsymbol{\theta}_{k}$. For large problems, computing the Jacobian matrix and its spectra this many times can become computationally problematic.
3.5 Overview
We have presented three single step transformations, each aimed at stabilizing a LOO importance sampler by bringing the proposal distribution closer to the LOO target in a different sense. The KL divergence and variance descent transformations correspond to the first-step of a forward Euler discretized solver for the corresponding gradient flow equations. The log-likelihood descent transformation repels the posterior parameter distribution away from the targeted LOO observation, undoing part of its contribution to the full data posterior. While each transformation uses gradient information, their Jacobians are simple to approximate, requiring no computation of full Hessian matrices.
@@ 4 Examples
In this manuscript we consider the broad widely-used class of models that have a sigmoidal parameterization
@@
For these models, the steps for each of the three transformations take the form
@@
$Q_{i}^{\textrm{Var}}(\boldsymbol{\theta})$ $=(-1)^{y_{i}}\pi(\boldsymbol{\theta}|\mathcal{D})e^{2\mu_{i}(\boldsymbol{\theta})(1-2y_{i})}\nabla\mu_{i}$ (31)
$Q_{i}^{\textrm{LL}}(\boldsymbol{\theta})$ $=-\left[y_{i}\big{[}1-\sigma(\mu_{i}(\boldsymbol{\theta})\big{]}-(1-y_{i})\sigma(\mu_{i}(\boldsymbol{\theta}))\right]\nabla\mu_{i},$ (32)
@@
$\quad+\left[\nabla\log\pi(\boldsymbol{\theta}|\mathcal{D})+2(1-2y_{i})\nabla\mu_{i}\right](\nabla\mu_{i})^{\intercal}\Big{\}}$ (34)
$\mathbf{J}_{T_{i}^{\textrm{LL}}}(\boldsymbol{\theta})=\mathbf{I}+h\Big{\{}(1-\sigma(\mu_{i}(\boldsymbol{\theta}))\sigma(\mu_{i}(\boldsymbol{\theta}))\nabla\mu_{i}(\nabla\mu_{i})^{\intercal}$
$\quad-\left[y_{i}\big{[}1-\sigma(\mu_{i}(\boldsymbol{\theta}))\big{]}-(1-y_{i})\sigma(\mu_{i}(\boldsymbol{\theta}))\right]\nabla\nabla\mu_{i}\Big{\}},$ (35)
@@
and $\pi(\boldsymbol{\theta})$ is the prior.
Here we will consider two popular sub-families of sigmoidal models.
4.1 Logistic Regression (LR)
LR is a sigmoidal model where $\mu_{i}(\boldsymbol{\theta})=\mathbf{x}_{i}^{\intercal}{\boldsymbol{\beta}}$, So, $\nabla_{\boldsymbol{\beta}}\mu_{i}=\mathbf{x}_{i}$, and $\nabla\nabla\mu=\mathbf{0}$. Because the Hessian of $\mu$ vanishes, the Jacobian of the function $Q_{i}$ for each of the three functions is a rank-one matrix and has only a single non-zero eigenvalue. LR admits exact Jacobian determinants for each of the three transformations:
@@ 4.1 Logistic Regression (LR)
×$\quad\quad\quad\quad\quad\times\mathbf{x}_{i}^{\intercal}\left[\nabla\log\pi(\boldsymbol{\theta}|\mathcal{D})+2(1-2y_{i})\mathbf{x}_{i}\right]\Big{|},$ (38)
$\mathcal{J}_{T^{\textrm{LL}}_{i}}(\boldsymbol{\theta})=|1+h\mathbf{x}_{i}^{\intercal}\mathbf{x}_{i}\sigma(\mu_{i}(\boldsymbol{\theta}))(1-\sigma(\mu_{i}(\boldsymbol{\theta}))|.$ (39)
4.2 Bayesian (ReLU) Neural Networks
Bayesian ReLU-nets [Lee, 2000, Ghosh and Doshi-Velez, 2017, Choi et al., 2018, Kristiadi et al., 2020, Bhadra et al., 2019] are piecewise linear [Sudjianto et al., 2021, Wang, 2022, Montúfar et al., 2014, Sudjianto et al., 2020] extensions to regression models. Being locally linear, these models have block-sparse Hessians and are also amenable to some limited degree of interpretability [Sudjianto et al., 2020, Chang et al., 2023]. One may write an $L$-layer ReLU Bayesian neural network recursively
@@ 4.2 Bayesian (ReLU) Neural Networks
These models are governed by the equations $\mu=\mathbf{W}_{2}a(\mathbf{z}_{1})+b_{2}$ and $\mathbf{z}_{1}=\mathbf{W}_{1}\mathbf{x}$, where $\mathbf{W}_{2}\in\mathbb{R}^{1\times d},$ $b_{2}\in\mathbb{R}$, $\mathbf{W}_{1}\in\mathbb{R}^{d\times p},$ $\mathbf{b}_{1}\in\mathbb{R}^{d}.$ This model has the first order derivatives ${\partial_{({W}_{2})_{1i}}}\mu=a(z_{1})_{i}$, ${\partial_{({W}_{1})_{ij}}}\mu=(W_{2})_{1i}a^{\prime}((z_{1})_{i})x_{j}$, ${\partial_{b_{2}}}\mu=1.$ The only non-zero components of the Hessian matrix for $\mu$ are the mixed partial derivatives
@@
$\mathbf{v}_{k}^{\pm}=\left(\begin{matrix}\tilde{\mathbf{u}}_{k}/\sqrt{2|\mathbf{u}_{k}|^{2}}\\ \pm\mathbf{e}_{k}/\sqrt{2}\\ 0\end{matrix}\right),$ (46)

### v2

# Gradient-flow adaptive importance sampling for Bayesian leave one out cross-validation with application to sigmoidal classification models

@@ Abstract
We introduce gradient-flow-guided adaptive importance sampling (IS) transformations for stabilizing Monte-Carlo approximations of leave-one-out (LOO) cross-validated predictions for Bayesian models. After defining two variational problems, we derive corresponding simple nonlinear transformations that utilize gradient information to shift a model’s pre-trained full-data posterior closer to the target LOO posterior predictive distributions. In doing so, the transformations stabilize importance weights. The resulting Monte Carlo integrals depend on Jacobian determinants with respect to the model Hessian. We derive closed-form exact formulae for these Jacobian determinants in the cases of logistic regression and shallow ReLU-activated artificial neural networks, and provide a simple approximation that sidesteps the need to compute full Hessian matrices and their spectra. We test the methodology on an $n\ll p$ dataset that is known to produce unstable LOO IS weights.
@@ 1 Introduction
In Bayesian inference, multiple models are often fit to data and then a selection procedure is applied to decide which model is best. A commonly used criterion for selection is how well a model accurately predicts out-of-sample data, which automatically optimizes fit while avoiding over-fitting. Prediction accuracy is most naturally estimated using cross-validation (CV). A model is trained with some of the data and tested against the unused data. However, estimates of out-of-sample model metrics using train-test splitting are noisy [Dietterich, 1998, Kohavi, 1995] unless computationally expensive k-fold cross-validation (i.e. fitting the model multiple times on cross-over the entire dataset) is employed [Rodriguez et al., 2010, Wong and Yeh, 2020].
Although $N$-fold, also known as leave one out (LOO), CV is the most expensive, there exist computationally efficient LOO techniques that completely avoid refitting. For example, the Akaike Information criteria (AIC) and Bayesian variants [Stone, 1977, Watanabe, 2010, Gelman et al., 2014, Watanabe, 2013] are asymptotic approximations of LOO CV. For Bayesian models, a more precise way to perform LOO CV is to use importance sampling [Vehtari et al., 2017, Piironen and Vehtari, 2017a], which works by using the full data posterior measure as a proposal distribution for each data point’s LOO posterior measure. These methods preclude refitting the model by approximating each LOO posterior measure using full data posterior measure. However, in cases where the LOO measure and full measure are very different, importance sampling can fail [Piironen and Vehtari, 2017a]. To ameliorate this possibility, we introduce an adaptive importance sampling method for LOO CV based on using transformations that bring the proposal distribution closer to LOO posteriors. We derive these transformations by defining gradient flows that minimize given statistical objective. While the transformations are model-dependent, the method is made completely general when using autograd for computing model gradients – for computational efficiency we derive the transformations exactly for a large class of classification models.
@@ 2 Preliminaries
Notation: We denote vectors (assumed to be column vectors unless otherwise stated) using bold-faced lowercase symbols, and matrices using bold-faced uppercase symbols. Given a matrix $\mathbf{W}=(w_{ij})$, the $i$-th row is denoted $\mathbf{w}_{i,}$ and $j$-th column is denoted $\mathbf{w}_{,j}$.
@@
Importance sampling-based approximate leave one out cross validation (IS-LOO): Suppose that one has pre-trained a Bayesian model such that one is able to sample its posterior parameters
$\boldsymbol{\theta}_{s}\overset{\mathrm{iid}}{\sim}\pi(\boldsymbol{\theta}|\mathcal{D}).$ Our objective is to use knowledge of this full-data posterior distribution to estimate how the model would behave if any single point is left out at training. One can relate the full-data model to the model with observation $i$ left out using the Bayesian update equation
@@
which is a Fredholm integral equation of thge with respect second-kind to $\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})$. This integral equation is in-practice difficult to solve due to the typically-high dimensionality of $\boldsymbol{\theta}.$
Rather than directly inverting Eq. 1, our starting point is the observation that Eq. 1 implies
@@
which provides the ratio of densities between a distribution we know (the full-data posterior $\pi(\boldsymbol{\theta}|\mathcal{D})$) and a distribution whose statistics we would like to compute (the point-wise LOO posterior $\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})$). To use the former to compute statistics of the latter we turn to Monte-Carlo [Barbu and Zhu, 2020, Robert and Casella, 2013] - the use of statistical sampling to compute a desired quantity (typically an integral). Importance Sampling (IS) is a Monte-Carlo method where one computes expectations with respect to a target distribution by taking a weighted average of samples with respect to a given proposal distribution. For an integrable function $f$,
@@
$=\int f(\boldsymbol{\theta})\frac{\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})}{\pi(\boldsymbol{\theta}|\mathcal{D})}\pi(\boldsymbol{\theta}|\mathcal{D})\mathrm{d}\boldsymbol{\theta}=\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}}\left[f(\boldsymbol{\theta})\nu_{i}(\boldsymbol{\theta})\right].$ (3)
We approximate Eq. 3 by sampling over $\boldsymbol{\theta}_{k}\overset{\mathrm{iid}}{\sim}\pi(\boldsymbol{\theta}|\mathcal{D})$, and computing the Monte-Carlo integral
@@
LOO-based metrics: A Bayesian analogue to the Aikaike Information Criterion (AIC) can be computed using the LOO information criterion (LOO-IC):
@@
Weight stabilization: Often it is the case that using the computed posterior $\pi(\boldsymbol{\theta}|\mathcal{D})$ as the proposal distribution for importance sampling has slow convergence properties – the $1/\ell$ importance weights are known to have large or unbounded variance [Peruggia, 1997], making the importance sampler estimate for LOO noisy.
Two practical model agnostic methods for controlling the tail of the importance weights are through weight truncation [Ionides, 2008] and Pareto smoothing  [Vehtari et al., 2024, 2017]. Pareto smoothing replaces the largest $M$ weights with their corresponding rank-values from a fitted generalized Pareto-distribution [Zhang and Stephens, 2009]. Pareto smoothed importance sample (PSIS)-based LOO implementations are widely available in software packages such as Stan and ArviZ. However, PSIS-LOO is insufficient in problems where the Pareto distribution does not well-fit the tail distribution of importance weights – where the estimated Pareto shape parameter $\hat{k}$ exceeds $0.7$ – this parameter is a useful criterion for accessing the validity of a given importance sampler. In such a case, it would be desirable to perform an additional model-specific controlled transformation on the proposal distribution that will induce more efficient computations.
@@ 2.1 Adaptive importance sampling
We apply the concept of adaptive importance sampling [Bugallo et al., 2017, Cornuet et al., 2011, Elvira and Martino, 2022] to transform the posterior distribution to be closer to the LOO distribution $\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})$ (relationships between the different distributions is depicted in Fig. 1).
@@
Using variational posteriors: For computational expediency, variational methods are often used in place of MCMC for Bayesian inference, obtaining a variational approximation $\hat{\pi}(\boldsymbol{\theta}|\mathcal{D})$ to the true posterior, where $\hat{\pi}$ lies within a given family of transformed probability distributions. In problems where one expects a substantial discrepancy between the true posterior and $\hat{\pi},$ one may correct for this discrepancy by noting that
@@ 3 Methods
Our focus is on classification models where a vector of covariates $\mathbf{x}\in\mathbb{R}^{p}$ is used to estimate the probability of an outcome labeled by $y\in\{0,1\}$ with likelihood function $\ell$:
where $\pi(\boldsymbol{\theta})$ is the joint prior distribution over the parameter vector $\boldsymbol{\theta}$, corresponding to the likelihood function $\ell(\boldsymbol{\theta})$. There are usually many plausible choices for $\ell,$ even within the same family of models (for example different sets of features in logistic regression) – one needs a method to evaluate different models to inform decisions such as model selection.
Eq. 11 is valid for an arbitrary bijection $T_{i}.$ The objective of using transformations is to shift the proposal distribution closer to the targeted LOO distribution for each observation – to partially invert the Bayes rule. To this end, we motivate two different transformations of the form
@@
3.1 Gradient flow transformations
@@
Starting from the associated variational problem (Appendix S.1.2), and applying the same rationale that went into developing the KL-descending transformation, one arrives at the single-step variance-reducing transformation,
@@
Resolving the posterior density: Both the KL (Eq. 21) and variance (Eq. 22) descent transformations take steps proportional to the posterior density $\pi(\boldsymbol{\theta}|\mathcal{D}).$ If a variational approximation for $\pi({\boldsymbol{\theta}}|\mathcal{D})$ is available, using it in Eqs. 21 and 22 as a stand-in for the posterior density helps simplify the computation of the transformations and their Jacobians, particularly when using mean-field or low-order Automatic Differentiation Variational Inference (ADVI) [Kucukelbir et al., 2017, Blei et al., 2017].
In the absence of variational approximation, one may evaluate the posterior densities exactly using the Bayes rule, absorbing the unknown normalization constant $\mathcal{Z}$ into the step size $h.$ Then, the KL and variance descent transformations are
@@
and $T^{\textrm{Var}}_{i}(\boldsymbol{\theta})=\boldsymbol{\theta}+hQ_{i}^{\textrm{Var}}(\boldsymbol{\theta}),$ where
@@
Step size selection The KL-divergence and variance descent transformations correspond to a forward Euler solver on the respective gradient flow equations. According to linear stability analysis, Euler’s method has the conditional stability criteria $h<{2}/{\max_{k}|\mathrm{Re}(\lambda_{k})|}$ where $\lambda_{k}$ are the eigenvalues of the Jacobian of the system (Jacobians of the functions $Q_{i}$). In each case the structure of the Jacobian admits cheap approximations of $\lambda_{k}$. However, for nonlinear systems, this criterion is not sufficient for achieving stability.
@@
$h_{i}=\rho\min_{s,\alpha}\left\{\left|{\sqrt{\Sigma_{\alpha,\alpha}}}/{Q_{i}(\boldsymbol{\theta}_{s})_{\alpha}}\right|\right\}$ (26)
@@
Jacobian determinant approximation: For either single-step transformations, one may approximate $\left|\mathbf{J}_{T_{i}}\right|$ by noting that
@@
and truncating to $\mathcal{O}(h),$ sidestepping the computation of Hessian matrices and their spectra. Note that any higher order terms in this expansion require characterization of the spectra of $\nabla Q_{i},$ for each observation $i$, and for each sampled parameter $\boldsymbol{\theta}_{k}$. For large problems, computing the Jacobian matrix and its spectra many times can become computationally problematic.
Overview: We have presented two gradient-flow descending transformations, each aimed at stabilizing a LOO importance sampler by bringing the proposal distribution closer to the LOO target in a different sense. The KL divergence and variance descent transformations correspond to the first-step of a forward Euler discretized solver for the corresponding gradient flow equations. The log-likelihood descent transformation repels the posterior parameter distribution away from the targeted LOO observation, undoing part of its contribution to the full data posterior. While each transformation uses gradient information, their Jacobians are simple to approximate, requiring no computation of full Hessian matrices.
@@ 4 Examples
In this manuscript, we consider the broad widely-used class of models that have a sigmoidal parameterization
@@
For these models, the transformations take the form
@@
$Q_{i}^{\textrm{Var}}(\boldsymbol{\theta})$ $=(-1)^{y_{i}}\pi(\boldsymbol{\theta}|\mathcal{D})e^{2\mu_{i}(\boldsymbol{\theta})(1-2y_{i})}\nabla\mu_{i},$ (30)
@@
$\quad+\left[\nabla\log\pi(\boldsymbol{\theta}|\mathcal{D})+2(1-2y_{i})\nabla\mu_{i}\right](\nabla\mu_{i})^{\intercal}\Big{\}},$ (33)
@@
and $\pi(\boldsymbol{\theta})$ is the prior. Here we will consider two popular sub-families of sigmoidal models.
Logistic Regression (LR): LR is a sigmoidal model where $\mu_{i}(\boldsymbol{\theta})=\mathbf{x}_{i}^{\intercal}{\boldsymbol{\beta}}$, So, $\nabla_{\boldsymbol{\beta}}\mu_{i}=\mathbf{x}_{i}$, and $\nabla\nabla\mu=\mathbf{0}$. Because the Hessian of $\mu$ vanishes, the Jacobian of the function $Q_{i}$ for each of the functions is a rank-one matrix and has only a single non-zero eigenvalue. LR admits exact Jacobian determinants for each of the transformations:
@@
×$\quad\quad\quad\quad\quad\times\mathbf{x}_{i}^{\intercal}\left[\nabla\log\pi(\boldsymbol{\theta}|\mathcal{D})+2(1-2y_{i})\mathbf{x}_{i}\right]\Big{|}.$ (36)
Bayesian (ReLU) Neural Networks: Bayesian ReLU-nets [Lee, 2000, Ghosh and Doshi-Velez, 2017, Choi et al., 2018, Kristiadi et al., 2020, Bhadra et al., 2019] are piecewise linear [Sudjianto et al., 2021, Wang, 2022, Montúfar et al., 2014, Sudjianto et al., 2020] extensions to regression models. Being locally linear, these models have block-sparse Hessians and are also amenable to some limited degree of interpretability [Sudjianto et al., 2020, Chang et al., 2023]. One may write an $L$-layer ReLU Bayesian neural network recursively
@@
These models are governed by the equations $\mu=\mathbf{W}_{2}a(\mathbf{z}_{1})+b_{2}$ and $\mathbf{z}_{1}=\mathbf{W}_{1}\mathbf{x}$, where $\mathbf{W}_{2}\in\mathbb{R}^{1\times d},$ $b_{2}\in\mathbb{R}$, $\mathbf{W}_{1}\in\mathbb{R}^{d\times p},$ $\mathbf{b}_{1}\in\mathbb{R}^{d}.$ This model has the first-order derivatives ${\partial_{({W}_{2})_{1i}}}\mu=a(z_{1})_{i}$, ${\partial_{({W}_{1})_{ij}}}\mu=(W_{2})_{1i}a^{\prime}((z_{1})_{i})x_{j}$, ${\partial_{b_{2}}}\mu=1.$ The only non-zero components of the Hessian matrix for $\mu$ are the mixed partial derivatives
@@
$\mathbf{v}_{k}^{\pm}=\left(\begin{matrix}\tilde{\mathbf{u}}_{k}/\sqrt{2|\mathbf{u}_{k}|^{2}}&\pm\mathbf{e}_{k}/\sqrt{2}&0\end{matrix}\right)^{\intercal},$ (43)